Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 629     Entry time: 16 Jul 2025, 09:51
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 16.07.2025 
General Announcement:

- Martin Doering informs on the LIMP-25 on cw30: 
1. On PLCs only the GATT will have an update next week; 
2. Ramp-up phase will start next week with tunnel racks, balcony rooms and Lasers; 
3. XHQ rack room doors will remain open next week - please do not enter rooms; 
4. SA2 HXS electrical work will finish this week; 
5. XHEXP1 UPS extension work is postponed and will not happen on 22-23.07. Once new date is available it will be communicated; 
6. Electrical work in XHQ: lab areas (level 0) being affected until 21.07; 
7. Electrical work on the XHQ offices of level 1 starts on 22.07 until 25.07; 
8. XHQ level 2 office areas are affected with electrical work from 29.07 to 01.08. 

- Martin Doering updates on the cw29 (current week): 
1. Today from 10AM to 11:30 craning activities are planned for HXS area; Electrical maintenance in E0/lab area are also affected on Thursday in the same way as today.

- Michael Prollius informs on the fire safety for SA2 experimental area: the smoke detection system is currently OFF due to the electrical work being done. 

CTRL:
- Nothing to report; Karabo to be ramped up next week from 22.07 to 25.07.

DA:
- Nothing to report.

EEE:
- Nothing major; temperatures of SA1 balcony rooms raised more than other areas - to be monitored to see if PLCs shall also be switched off.

VAC:
- Not present.

XPD:
- Nothing to report.

DET:
- Nothing to report.

ITDM: 
- Balcony rooms being monitoring temperatures.
- Doors to the rooms are opened and there is a trend of slow raise of temperature ongoing and flow of air from the Hall to be improved with fans.

XRO:
- Nothing to report; activities are progressing as planned;

LAS:
- Progressing with tasks.

UNSYS:
- Nothing to be reported.

SPB:
- Nothing to report, all good.

FXE:
- Nothing to report.

MID:
- Nothing to report.

HED:
- Ongoing work depends on the return of electricity.

SCS:
- Emergency OFF tests were delayed and Tobias C. informed it will happen today. 

SQS:
- Nothing to report.

SXP:
- Not present.
ELOG V3.1.4-7c3fd00